OpenStudy (anonymous):

use the quadratic formula to solve for exact values for x x - (1/x) = 1

OpenStudy (anonymous):

\[x^2 - x -1=0 \]\[x = \frac{1 \pm \sqrt{1 + 4}}{2}\]